OpenStudy (anonymous):

A trapezoid has bases with lengths of 15 cm and 19 cm and an altitude with a length of 12 cm. What is the area of the trapezoid? A. 136 cm2 B. 204 cm2 C. 408 cm2 D. 816 cm2

OpenStudy (anonymous):

area of trap = altitude(average of bases) = a(b1+b2)/2 = 12(15+19)/2= 6*34 = 204